SRP Telecons. The telecons themselves went smoothly, but there were some scoring issues. Several reviewers (including chairs) reversed their scores (higher numbers being a better proposal). It was not trivial to fix these in the PST, especially for the first case, and in the past we have always caught these before the reviews were normalized. There was some discussion about what to do, if anything: remind reviewers of the scoring rules; generate statistic to flag potential problems; or since most scoring systems use higher numbers as better (e.g., NSF, NASA, Arecibo) change NRAO to conform.
NRAO PHT Status. The PHT updates have been released to production. Per the suggestion by BarryClark we agreed that EVL-3202 (TAC csv file) was sufficient to use for 16B even though it still had some issues. RickLively needs to fix up the joint external Chandra-NRAO proposals since the software was not updated to account for the fix we made to these in the PST.
GB PHT Status. ToneyMinter reported before the meeting that there were no issues. The software folks will exercise the import/export functionality.
Large Proposal Telecon. The list of Large proposals have been sent to the TAC along with a Doodle poll to find a time. Because of the delay in releasing the normalized linear scores the GBT TAC report will not be ready for the telecon. Given that there are not too many GBT proposals, and only two Large proposals, we will defer any scheduling issues to the f2f meeting.
TAC f2f Meeting. All logistical and technical requests have been sent to DavidFrayer, the local contact person, and no issues were raised. At least one chair will not be able to attend.